what is 2a squared if a =4 is it 256 or 32
Answer by rothauserc(4718)

(2a)^2 = 2^2 * a^2 = 4*a^2, now if a = 4 we have
4*4^2 = 4*16 = 64
another way to look at this is
(2*4)^2 = 8^2 = 64